### v1

@@ 3.3 Cache Prior Reranking
To address this, we use a cache prior that directly manipulates the router logits $\mathbf{z}=G(\mathbf{x})$ to increase the probability of selecting experts already present in cache. Importantly, we use the manipulated logits $\mathbf{z}^{\prime}$ only to find a new ranking vector $\mathbf{r}^{\prime}$ as defined in Equation 2. We still use the unmodified router logits to compute the expert weights used in Equation 1. Figure 3 provides an overview of our proposed cache-aware routing method. Let the bitmask $\mathbf{m}_{t}\in\{0,1\}^{N}$ denote the state of the cache following the generation of the $(t-1)$-th token.
As with our other methods, we aim to ensure that the Top-J experts are always selected, regardless of wether they are in the cache. We achieve this by adding them to the bitmask $\mathbf{m}_{t}$, resulting in an updated bitmask denoted as $\tilde{\mathbf{m}}_{t}$. We then boost the logits corresponding to the bitmask as follows:
@@
where $\lambda\in[0,1]$ is a scaling factor that determines the influence of the cache state on the logits. Here, $\Delta(\mathbf{z})$ is defined as:
@@
which represents the range in logits for this layer that we estimate using a running average over datapoints and over tokens.
In essence, this method involves adding a fraction of $\Delta(\mathbf{z})$ to the logits of experts that are already in the cache. By utilizing the true range of weights, our cache prior becomes adaptive, allowing a single hyperparameter to yield distinct effects across different layers and tokens.
We provide an ablation on the choice of $\Delta$ in Section 4.6. Additionally, we explore the possibility of directly learning $\Delta$ in Appendix, Section D.
@@ 4 Experiments
Model Params Experts
Total Active Shared Total Top-k Params
Mixtral-8x7B 46.7B 13B 8 2 176M
Phi-3.5-MoE 41.9B 6.6B 16 2 79M
DeepSeek-V2 15.92B 2.8B 2 (64+2) (6+2) 8.6M
Qwen1.5-MoE 14.3B 2.7B 4 (60+4) (4+4) 8.6M
Table 1: The MoE architectures used in our experiments. Column Params in Experts indicates the amount of parameters per expert.
@@
Figure 5: MMLU (5 shots) task evaluation.
Figure 6: GSM8K (8 shots) task evaluation.
@@
In this section, we evaluate the effectiveness of our cache-aware routing method (dubbed Prior for simplicity) and compare it to our Pruning, Max-Rank, and Cumulative Sum Thresholding baselines described in Section 3. We implement all methods on four SoTA MoE models: DeepSeek-V2-Lite DeepSeek-AI et al. (2024), Qwen1.5-MoE-A2.7B Qwen Team (2024), Phi-3.5-MoE Abdin et al. (2024), and Mixtral-8x7B Jiang et al. (2024). The architectural details of these models are summarized in Table 1.
The MoE layers in DeepSeek-V2-Lite include 2 shared experts activated for every token and 64 selectable experts, out of which 6 are activated per token, resulting in 2.4B active parameters per token out of a total of 15.7B parameters. Similarly, Qwen1.5-MoE-A2.7B has 4 shared experts and 60 selectable experts, with 4 activated for each token, leading to 2.7B active parameters (total 14.3B parameters). Phi-3.5-MoE comprises 16 experts using Top-2 routing, resulting in 6.6B active parameters (total 42B parameters). Lastly, Mixtral-8x7B includes 8 experts with Top-2 routing, resulting in 12.9B active parameters (total 46.7B parameters).
To assess the performance of our method, we conduct experiments three tasks: language modeling, Massive Multitask Language Understanding (MMLU), and Grade School Math (GSM8K). For language modeling, we report perplexity and cache miss rate on the validation subset of the WikiText-2-raw-v1 dataset Merity et al. (2017). Next, we evaluate downstream performance on the MMLU dataset Hendrycks et al. (2021) and the GSM8K dataset Cobbe et al. (2021). The MMLU dataset consists of multiple-choice questions spanning 57 subjects, while GSM8K evaluates multi-step reasoning on math problems. For both datasets, we report accuracy and cache miss rate.
In our experiments, we examine trade-offs between model accuracy and cache miss rate under varying DRAM cache size limits. For each method, we sweep over hyperparameters to create a pareto front of performance vs. cache miss rate.
In addition to the baseline results, each plot includes the performance of the model utilizing the Least Recently Used (LRU) caching policy. All of our routing methods implement an LRU cache, but the LRU baseline specifically refers to the model that employs the LRU cache without any re-ranking strategy. As a result, this baseline maintains accuracy without any trade-offs. In contrast, our other methods incorporate routing strategies that balance accuracy with LRU friendliness, allowing for improved cache utilization while still aiming to optimize performance.
@@ 4.1 Implementation Details
Keeping Top-J Experts
As observed in Section 2.3, certain Top-J experts are crucial for model performance. To prevent these experts from being overwritten by others in the cache, we always select them, even if they are not in the cache. In the cases of Phi and Mixtral, where only two experts are activated, we use $J=1$. For Qwen and DeepSeek, where the first major perplexity drop occurs at $J=2$, we select $J=2$.
We demonstrate the impact of the specific choice of Top-J in Figure 4. As shown, retaining the Top-J experts is crucial for the performance of baseline methods. Simultaneously, impact to our cache-aware routing method is less clear. However, its effect on our cache-aware routing method is less pronounced. To ensure a fair comparison and simplify the approach, we use the same Top-J configuration across all methods. Further discussion is presented in Section 4.6.
Dataset-Specific Details
@@
Pruning
In our pruning baseline, we select a rank $h\leq k$ and discard all experts with a rank greater than or equal to $h$. Note that we still use $k$ to normalize the cache miss rate, as indicated in Equation 4.
LRU
For all methods in this work, when the cache is full, we rely on the LRU eviction policy to drop experts and free up space. Unlike traditional paging, there is no inherent order for incoming elements in each MoE layer, as an expert’s importance for the current token does not necessarily indicate its future relevance. We choose to add elements to the LRU stack based on their router weights, prioritizing experts with the highest weights for addition to (and subsequent eviction from) the cache. We observed that this approach yielded slightly better performance compared to the inverse method, as shown in Figure 10 in Appendix C.
@@ 4.2 Language Modeling Evaluation
In this section, we assess the effect of our proposed cache-aware routing method on the language modeling capabilities of MoE models. Figure 4 presents trade-off curves for perplexity versus cache miss rate of our Prior approach compared to the Pruning, Max-Rank, and Cumsum-threshold methods with a cache size equal to half the number of experts. The results show that our approach consistently outperforms all baselines across all ranges of perplexity and cache miss rate. Among the baseline methods, Pruning performs the worst, indicating that cache-informed expert replacement is essential for maintaining accuracy. Max-Rank consistently surpasses Pruning, while Cumsum-Threshold consistently outperforms Max-Rank. Ultimately, our Prior method Pareto-dominates all other approaches.
We also evaluate language modeling performance with only a quarter of experts cached, as shown in Figure 12 in Appendix, Section B. Our method’s robustness to different cache sizes makes it appealing for deployment across a range of devices.
@@ 4.3 Downstream Tasks Evaluation
Figure 5 and Figure 6 present results on the MMLU and GSM8K benchmarks, respectively. Our Prior method’s Pareto curve consistently outperforms other methods, often showing significant cache miss rate reductions with no accuracy loss compared to baseline LRU. Increasing the cache bias $\lambda$ improves cache miss rates until a minimum is reached due to the Top-J selection guarantee. We also observe that GSM8K has noisier accuracy results, as can be expected for a generative task. Importantly, while accuracy may fluctuate, adjustments to the hyperparameters for each routing method —namely $M$, $p$, and $\lambda$— consistently leads to a predictable effect on cache miss rates.
Interestingly, we observe that improved cache consistency (reduced miss rate) can lead to increased accuracy on both benchmarks, potentially due to a regularization effect from consistent expert routing.
Figure 7 shows a qualitative visualization of cache hits/misses for a GSM8K sample, comparing LRU and Prior($\lambda$=0.5) methods, with cache miss rates of $\sim$23and $\sim$15%, respectively, for Phi-3.5-MoE-instruct. It is evident that the Prior method results in fewer cache misses, and the cache lifetimes for this method are longer than those of the LRU baseline. These observations are further supported by Table 2, which presents the average cache miss rate and cache lifetime for the Wikitext dataset.
Model Cache Size Strategy Lifetime Miss Rate (%)
Qwen1.5-MoE 30 / 60 LRU 22 35
Prior 111 7
DeepSeek-V2 32 / 64 LRU 20 28
Prior 188 3
Phi-3.5-MoE 8 / 16 LRU 22 22
Prior 55 9
Mixtral-8x7B 4 / 8 LRU 5 40
Prior 10 21
Table 2: Cache sizes, average cache lifetimes, and miss rates for LRU and prior caching strategies with $\lambda=0.5$ on the Wikitext validation set.
@@ 4.4 On-Device Deployment
To evaluate the effectiveness of our cache-aware routing technique in real-world scenarios, we deployed the Qwen1.5-MoE-A2.7B model with our cache-aware routing on two mobile devices (with 12GB and 16GB RAM) equipped with some of the latest Qualcomm Snapdragon®  [^1] Snapdragon branded products are products of Qualcomm Technologies, Inc. and/or its subsidiaries. processors running Android 14. Using llama-cpp Gerganov (2023) for CPU-based deployment, we modified the implementation to add both the LRU caching for experts and our Prior algorithm. Additionally, we enabled memory locking (mlock) to prevent the Android OS from offloading expert weights from memory, which would otherwise require reloading for each token.
@@
In the first setting, we reserved 2GB of memory on the 12GB device, limiting available memory to 10GB (shared with the OS). Here, the cache size was limited to 45 experts per layer (out of 60). In the second setting, the 16GB device hosted the 8-bit quantized Qwen Model with a cache size limit of 30 experts per layer.
To determine optimal cache sizes for each setting, we initially deployed the model on both devices using the LRU caching policy across various cache sizes. Results, shown in Figure 9, indicate that exceeding a certain cache size reduces available memory, causing the OS to reload uncached components (e.g., kv-cache, activations) for each token, which significantly slows down inference. We used the highest LRU throughput as a reference point (1x) to demonstrate the speedup from our method independently of wall clock variations due to implementation specifics.
@@ 4.5 Temporal Consistency of Expert Selections
In Table 2, we assess the impact of our method on average cache lifetime, defined as the average number of time steps (tokens generated) an expert remains in memory before being offloaded. Higher cache lifetime indicates better memory efficiency. We calculated this measure over the entire WikiText validation set, averaged across tokens.
Our approach significantly extends cache lifetimes. For example, using LRU, each expert in both Qwen and DeepSeek models stays in memory for only  22 tokens, often less than a few seconds on modern devices. With our method, experts remain in memory for 59×5-9$\times$5 - 9 × longer, reducing the frequency of weight loading from Flash and improving throughput.

### v2

@@ 3.3 Cache-Prior Reranking
To address this, we use a Cache-Prior that directly manipulates the router logits $\mathbf{z}=G(\mathbf{x})$ to increase the probability of selecting experts already present in cache. Importantly, we use the manipulated logits $\mathbf{z}^{\prime}$ only to find a new ranking vector $\mathbf{r}^{\prime}$ as defined in Equation 2. We still use the unmodified router logits to compute the expert weights used in Equation 1.
Figure 3 provides an overview of our proposed cache-aware routing method. Let the bitmask $\mathbf{m}_{t}\in\{0,1\}^{N}$ represent the state of the cache following the generation of the $(t-1)$-th token, where each bit indicates whether an expert is in the cache or not.
To ensure the top-$J$ experts are always selected, regardless of their presence in the cache, we can optionally add them to the bitmask $\mathbf{m}_{t}$. This results in an updated bitmask, $\tilde{\mathbf{m}}_{t}$. Using this updated bitmask, we then boost the logits for the corresponding experts as follows:
@@
where $\lambda\in[0,1]$ is a scaling factor that determines the influence of the cache state on the logits. This scaling factor allows our method to interpolate between the original routing ($\lambda=0$, equivalent to baseline cache hit rate) and a fully cache-driven selection ($\lambda=1$, potentially zero cache misses, but higher perplexity). $\Delta_{\text{avg}}$ is defined as:
@@
which represents the range in logits for this layer that we estimate using a running average over sequences and tokens. In essence, our Cache-Prior method involves adding a fraction of the average logit range to the experts that are already in the cache. By utilizing the true range of weights, our Cache-Prior becomes adaptive, allowing a single hyperparameter to yield distinct effects across different layers and tokens. We provide an ablation on the choice of $\Delta$, including approaches for predicting it directly for each token, in Appendix D and E.
Figure 3: Our proposed Cache-Prior routing method adds a bias to the logits only for in-cache experts $\mathbf{m}_{t}$, encouraging their selection. The magnitude of the bias is determined by the average logit range, $\Delta_{\text{avg}}$, and the tradeoff parameter $\lambda$. The updated logits, $\mathbf{z}_{t}^{\prime}$, are used only for re-ranking experts, while the expert weights, $\mathbf{w}_{t}$, remain unchanged.
@@ 4 Experiments
Model Params Experts Footprint (int4)
Active Total Shared Total Top-k Params min max
Mixtral-8x7B 13B 46.7B 8 2 176M 6,5 GB - 23,4 GB
Phi-3.5-MoE 6.6B 41.9B 16 2 79M 3,3 GB - 21,0 GB
DeepSeek-V2 2.8B 15.9B 2 (64+2) (6+2) 8.6M 1,4 GB - 8,0 GB
Qwen1.5-MoE 2.7B 14.3B 4 (60+4) (4+4) 8.6M 1,4 GB - 7,2 GB
Table 1: The MoE architectures used in our experiments. Column “Experts” / “Params” indicates the number of parameters per expert. Column “Footprint” represents the total size of all model parameters and the expert cache under int4 quantization, with the cache size bounded by $k$ (minimum) and $N$ (maximum).
@@
Figure 5: The trade-off between MMLU (5 shots) task accuracy and cache miss rate. For each method, points along the curve form the Pareto front, showcasing the best achievable accuracy for a given cache miss rate.
Figure 6: The trade-off between GSM8K (8 shots) task accuracy and cache miss rate. For each method, points along the curve form the Pareto front, showcasing the best achievable accuracy for a given cache miss rate.
@@
In this section, we evaluate the effectiveness of our Cache-Prior routing method and compare it to out Pruning, Max-Rank, and Cumulative Sum Thresholding baselines, as described in Section 3. We implement all methods on four state-of-the-art MoE models: DeepSeek-V2-Lite (DeepSeek-AI et al., 2024), Qwen1.5-MoE-A2.7B (Qwen Team, 2024), Phi-3.5-MoE (Abdin et al., 2024), and Mixtral-8x7B (Jiang et al., 2024), with architectural details summarized in Table 1.
4.1 Experimental Setup
We conduct experiments on three tasks: language modeling (using the WikiText-2-raw-v1 dataset), MMLU, and GSM8K. For WikiText, we report perplexity and cache miss rate, while for MMLU and GSM8K, we report accuracy and cache miss rate. The MMLU dataset consists of multiple-choice questions across 57 subjects, and GSM8K evaluates multi-step reasoning for math problems. For all experiments, the cache miss rate is computed using the Least Recently Used (LRU) eviction policy, unless stated otherwise in ablations.
To assess the trade-offs between model performance and cache miss rate, we vary DRAM cache size limits and sweep hyperparameters for each method to generate Pareto fronts. All reported results in the trade-off plots are point estimates, obtained from a single pass over the entire dataset. As our algorithm and baselines are deterministic, repeated runs yield identical results. Each plot also includes the performance of the “Original Routing” baseline, where no re-ranking strategy is applied, maintaining accuracy without trade-offs. In contrast, our cache-aware routing methods balance accuracy with cache utilization.
@@ 4.2 Implementation Details
All routing strategies in this work use the LRU eviction policy for expert removal, where we impose an eviction order by removing experts with higher router weights first, as MoE layers select top-$K$ experts in parallel. For the pruning baseline, experts ranked $\geq h$ (with $1<h\leq k$) are discarded, and the cache miss rate is normalized using $k$ for a fair comparison with other methods.
Each cache-aware routing strategy has a hyperparameter to balance cache miss rate and task performance. We use the following values to generate Pareto curves: Pruning and Max-Rank use ${0,1,...,K}$, while Cumulative Sum Thresholding and Cache-Prior use 50 equidistant points in $[0,1]$. For guaranteed top-$J$ loading, we set $J=1$ for Mixtral and Phi-MoE models and $J=2$ for the granular Qwen-MoE and DeepSeek-MoE architectures.
@@ 4.3 Language Modeling Evaluation
We start with assessing the effect of our proposed cache-aware routing method on the language modeling capabilities of MoE models. Figure 4 presents trade-off curves for perplexity versus cache miss rate of our Cache-Prior approach compared to the Pruning, Max-Rank, and Cumsum-threshold methods with a cache size equal to half the number of experts. The results show that our approach consistently outperforms all baselines across all ranges of perplexity and cache miss rate. Among the baseline methods, Pruning performs the worst, indicating that cache-informed expert replacement is essential for maintaining accuracy. Max-Rank consistently surpasses Pruning, while Cumsum-Threshold consistently outperforms Max-Rank. Ultimately, our Cache-Prior method Pareto-dominates all other approaches.
We also evaluate language modeling performance with only a quarter of experts cached, as shown in Figure 15 in Appendix D. Our method’s robustness to different cache sizes makes it appealing for deployment across a range of devices.
@@ 4.4 Downstream Tasks Evaluation
Figure 5 and Figure 6 present results on the MMLU and GSM8K benchmarks, respectively. Our Cache-Prior method’s Pareto curve consistently outperforms other methods, often showing significant cache miss rate reductions with no accuracy loss compared to original routing. Increasing the cache bias $\lambda$ improves cache miss rates until a minimum is reached due to the top-$J$ selection guarantee. We also observe that GSM8K has noisier accuracy results, as can be expected for a generative task. Importantly, while accuracy may fluctuate, adjustments to the hyperparameters for each routing method —namely $M$, $p$, and $\lambda$— consistently leads to a predictable effect on cache miss rates.
Interestingly, we observe that improved cache consistency (reduced miss rate) can lead to increased accuracy on both benchmarks. We hypothesize this may stem from an implicit regularization effect, where greater temporal consistency in expert selection stabilizes predictions, although other factors likely contribute.
@@ 4.5 On-Device Deployment
To evaluate the effectiveness of our cache-aware routing technique in real-world scenarios, we deployed the Qwen1.5-MoE-A2.7B model with our cache-aware routing on two mobile devices (12GB and 16GB RAM) equipped with Qualcomm Snapdragon®footnotetext: Snapdragon branded products are products of Qualcomm Technologies, Inc. and/or its subsidiaries. processors running Android 14. Using llama-cpp (Gerganov, 2023) for CPU-based deployment, we modified the implementation to add both LRU caching for experts and our Cache-Prior algorithm. Additionally, we enabled memory locking (mlock) to prevent the Android OS from offloading expert weights from memory.
@@
In the first setting, we reserved 2GB of memory on the 12GB device, limiting available memory to 10GB (shared with the OS). Here, the cache size was limited to 30 experts per layer (out of 60). In the second setting, the 16GB device hosted the 8-bit quantized Qwen Model with a cache size limit of 45 experts per layer. To determine these optimal cache sizes for each setting, we initially deployed the model on both devices using the LRU caching policy across various cache sizes. Results, shown in Figure 14, indicate that exceeding a certain cache size reduces available memory, causing the OS to offload uncached components (e.g., KV-cache, activations) for each token, requiring reloading them from flash which significantly slows down inference. We used the highest LRU throughput as a reference point (1x) to demonstrate the speedup from our method independently of wall clock variations due to implementation specifics.
@@
To quantify the relationship between cache hit rate and token generation throughput, we conducted an experiment, deploying the 4-bit quantized Qwen model on the 16GB device and performed a sweep over the Cache-Prior trade-off parameter $\lambda\in\{0.1,0.3,0.5,0.7,0.9\}$. A random subset of the MMLU dataset, comprising both short (40-60 tokens) and long (300-400 tokens) prompts, was used for these measurements.
Figure 8 (left) illustrates the relation between cache hit rate and relative token generation throughput for two cache configurations: 30 experts and 45 experts cached per layer (out of 60 experts). For each configuration, the throughput achieved with standard LRU caching and original routing (equivalent to $\lambda=0$ where no expert swapping occurs) serves as the baseline (relative throughput of 1.0). The results demonstrate a near-linear positive relationship: as $\lambda$ increases, the cache hit rate improves, leading to a corresponding increase in throughput. This observed trend is logical as cache-hit rate linearly correlates with the number of flash reads. As loading from flash is the major performance bottleneck, the number of flash reads should correlate roughly linearly with the throughput.
Figure 8: Impact of cache hit rate and prompt length on token generation throughput for the 4-bit quantized Qwen1.5-MoE-A2.7B model on the 16GB device. (Left) Relationship between cache hit rate and relative throughput (normalized to the LRU baseline of 1.0 for each cache size) for cache sizes of 30 and 45 experts (out of 60), across varying $\lambda$ values $0.1-0.9$. (Right) Influence of prompt length on relative throughput across varying $\lambda$ values, with a cache size of 45 experts.
Furthermore, Figure 8 (right) examines the influence of prompt length on throughput for a cache size of 45 experts. Across all values of $\lambda$, longer input prompts yield higher throughput (Similar trend is observed for the cache size of 30 in Appendix F). This is likely attributable to the amortization of initial overheads (such as model loading and initial cache fills) over a greater number of generated tokens, and potentially more stable expert usage patterns for longer sequences.
@@ 4.6 Temporal Consistency of Expert Selections
Figure 7 shows a qualitative visualization of cache hits/misses for a GSM8K sample (tokens 1000-1100), comparing original routing and Cache-Prior ($\lambda$=0.5) methods, with cache miss rates of $\sim$23and $\sim$15%, respectively, for Phi-3.5-MoE-instruct. It is evident that the Cache-Prior method results in fewer cache misses and longer cache lifetimes compared to the original routing baseline.
These observations are further supported by Table 10, which presents average cache miss rates and cache lifetimes (the average number of time steps an expert remains in memory) for the Wikitext dataset. Longer cache lifetimes indicate better memory efficiency.
Using the original routing, each expert in Qwen and DeepSeek models stays in memory for only  22 tokens. With our method, experts remain in memory for 59×5-9$\times$5 - 9 × longer, reducing weight loading from Flash and improving throughput. MoE models with fine-grained implementation (Ludziejewski et al., 2024)—where the hidden dimension of each expert is smaller than a standard feed-forward layer (e.g., Qwen-MoE and DeepSeek-MoE)—are more cache-friendly compared to conventional MoE architectures like Phi-MoE and Mixtral.
Model Cache Size Routing Lifetime Miss Rate
Qwen1.5-MoE 30 / 60 Original $26\,(\pm 31)$ 35%
Cache-Prior $58\,(\pm 67)$ 16%
DeepSeek-V2 32 / 64 Original $19\,(\pm 17)$ 28%
Cache-Prior $76\,(\pm 90)$ 7%
Phi-3.5-MoE 8 / 16 Original $22\,(\pm 31)$ 22%
Cache-Prior $55\,(\pm 94)$ 9%
Mixtral-8x7B 4 / 8 Original $5\,(\pm 4)$ 40%
Cache-Prior $10\,(\pm 10)$ 21%
Figure 9: Cache sizes (number of experts that fit in cache / total number of experts in model), average cache lifetimes (in number of tokens), and cache miss rates for LRU and prior caching strategies with $\lambda=0.5$ on the Wikitext validation set.
Figure 10: Tradeoff between language modeling performance and the number of parameters loaded from Flash memory
. “Optimal” represents the oracle cache eviction bound.
4.7 Impact of MoE Design Choices on Caching Performance
While retraining MoE models with varying design choices is beyond this paper’s scope, our experiments offer insights into how architectural choices in pre-trained MoEs affect deployment-time caching. The evaluated MoE models (Table 1) represent diverse architectures, differing in granularity, routing mechanisms, expansion rates, and the use of shared experts.
Granularity
Our results suggest that model granularity (Ludziejewski et al., 2024) influences cache efficiency, particularly when combined with our cache-prior reranking method. We observe that granular MoEs (Qwen-MoE, DeepSeek-MoE) tend to be more resilient to the approximate routing introduced by expert swapping compared to non-granular models like Mixtral. For instance, while Qwen-MoE and Mixtral exhibit comparable initial LRU cache miss rates (35% and 40%, respectively) in Figure 4, our cache-prior method halves Qwen-MoE’s miss rate with only a 0.5% perplexity increase. Achieving a similar miss rate reduction for Mixtral incurs a much higher perplexity penalty (2.9%). DeepSeek-MoE demonstrates even greater resilience, halving its baseline miss rate with a negligible perplexity increase ( 0.1%).
Expansion Rate
The expansion rate, defined as the ratio of activated expert parameters to the total number of expert parameters (Ludziejewski et al., 2024), also appears correlated with cache performance. We observe that models with lower expansion rates generally exhibit lower baseline cache miss rates. Specifically, Phi-MoE, Qwen-MoE, and DeepSeek-MoE (all with an expansion rate of 0.125) tend to show better cache characteristics under LRU compared to Mixtral (expansion rate of 0.25).
Top-K
We observed no clear link between top-$K$ and overall cache-friendliness (e.g., Mixtral vs. Phi, both $k=2$). However, models with smaller $k$ (Mixtral, Phi) are more sensitive to top-$1$ expert swaps and benefit more from guaranteeing top-$1$ expert loading (Figure 2 and Figure 4). In contrast, Qwen ($k=4$) and DeepSeek ($k=6$) show greater resilience to expert swapping, although their use of shared experts might also contribute to this behavior.